A manufacturer produces 3 models of TV sets. Model A requires 2 hours of electronics work, 2 hours of assembly work, and 1 hour of finishing work. Model B requires 1, 3, and 1 hours of electronics, assembly, and finishing work respectively. Model C requires 3, 2, and 2 hours of the same work. In one week, there are 100 hours available for electronics work, 100 hours for assembly, and 65 hours for finishing. How many models of each type should be produced each week if all available work hours are to be used?
Answer #1

2 Depwalk Tv modela Iv model B Tv modela 1 Total available hour 100 Electronic work 2 Assembly work 2 3 100 65 Finishing withsolving equation ® and 24 + y À 32 Box: 3y + 2z. 100 - 100 - 2 4 + z , z = Salve equation and ( try to eliminate variable a:Solve @40 >@ 0 -24 + Z 0. -a ( y - Rz) of 30.) -2. - Ry + 42 - 60. - 2y + + z 20. + 42.60. - - 32 : -60. G620 from @ -Ry + ROfoon © 2x + 10+, 3x R0 = 100. 34 = 100 -60 -10 2X = 30 x 30 is a (x. 15 units x 15 units , y = lounits, z= rounits) there fir